# Bayesian Theory of Consciousness as Exchangeable Emotion-Cognition Inference

### v2

@@ Abstract
This paper presents a unified theoretical framework in which consciousness emerges as a cycle-consistent, affectively anchored inference process, recursively structured by the interaction between emotion and cognition. Drawing from information theory, optimal transport, and the Bayesian brain hypothesis, we formalize emotion as a low-dimensional structural prior and cognition as a specificity-instantiating update mechanism. The resulting emotion–cognition cycle minimizes joint uncertainty by using emotionally valenced states to guide cognitive selection, while cognitive appraisals reciprocally regulate emotional states. This recursive inference loop provides a computational bridge to the challenging problem of consciousness, reframing subjective experience as the informational footprint of temporally extended, affect-modulated simulation.
We extend this framework by introducing the Exchangeable Integration Theory of Consciousness (EITC), in which temporally distributed emotional–cognitive episodes are modeled as conditionally exchangeable samples drawn from a latent affective self-model. This latent variable enables integration by supporting a coherent cause–and–effect structure with nonzero irreducibility, and differentiation by preserving the specificity of each episode under a shared generative prior. Crucially, we unify this structure with the Bayesian theory of consciousness through Rao-Blackwellized inference, which marginalizes latent self-structure while enabling context-sensitive belief updates. This mechanism bridges exchangeability with stability, coherence, and goal-directed simulation, capturing both local adaptability and global consistency. These dynamics are formalized via theorems grounded in De Finetti’s exchangeability principle, information integration theory, and KL-regularized optimal transport. Together, these contributions suggest that consciousness is not a static property but an evolving, distributed architecture, shaped by emotion, refined by cognition, stabilized by exchangeability, and unified through recursive inference over latent self-structure.
@@ IV-B Access Consciousness as Temporally-Extended Affect-Modulated Inference
While cycle-consistency captures the temporal coherence of consciousness, a deeper functional insight emerges when we recognize that this consistency is mediated by the recursive interaction between emotion and cognition, framing access consciousness not merely as sustained inference, but as affectively guided, cognitively refined simulation. Next, we propose a unifying theoretical framework in which access consciousness emerges from a temporally extended, recursive inference process modulated by the interaction between emotion and cognition. This view integrates concepts from active inference [10], optimal transport theory [61], and the information bottleneck method [36]. Consciousness is framed not as a static state, but as a dynamic, cycle-consistent alignment between perceptual context and cognitive content over time for accessing internal goals. We first construct temporally-extended inference by connecting with the theory of optimal transport alignment [61].
@@
Let $\pi(\Psi_{1:T}|\Phi^{*})$ denote a goal-conditioned distribution over context trajectories induced by an affective anchor $\Phi^{*}$, and let $q(\Psi_{1:T})$ denote the inferred distribution over contextual states. Define a ground cost function $c:\Psi\times\Psi\rightarrow\mathbb{R}_{\geq 0}$, typically $c(\psi,\psi^{\prime})=\|\psi-\psi^{\prime}\|^{2}$. The Optimal Transport cost between the two distributions is given by:
@@
Temporally Extended Inference via Optimal Transport. Let $\Phi^{*}$ represent a goal or affective anchor, and let $\pi(\Psi_{1:T}|\Phi^{*})$ denote a simulated trajectory of contexts consistent with that goal, a product of inverted inference. Let $q(\Phi_{1:T}|\Psi_{1:T})$ represent the actual inferred trajectory of content under incoming context, a result of forward inference. We define the conscious inference process as a recursive optimization that minimizes the sum of context-content uncertainty and transport mismatch between simulated and actual trajectories: $\min_{q}\left[\sum_{t=1}^{T}\gamma_{t}H(\Phi_{t}|\Psi_{t})+\lambda\,\mathrm{OT}\left(\pi(\Psi_{1:T}|\Phi^{*}),q(\Psi_{1:T})\right)\right]$, where $\gamma_{t}$ encodes affective precision, modulating attention and inference strength based on emotion, $H(\Phi_{t}|\Psi_{t})$ is the conditional entropy of content given context at time $t$, and $\mathrm{OT}(\cdot,\cdot)$ is an optimal transport cost (e.g., Wasserstein distance [62]) aligning goal-simulated and observed trajectories.
@@ IV-C Self-Consciousness as Exchangeability-Regularized Inverted Inference
Remark: Violation of this transport alignment leads to disintegration of conscious coherence, as seen in dissociation, confusion, or neurodegeneration. A specific example of violating transport alignment in neurodegeneration can be found in Sec. VI.
V Toward a Bayesian Theory of Consciousness with Exchangeability
@@
V-A Exchangeable Integration Theory of Consciousness (EITC)
@@
Remark: The above theorem integrates Bayesian inference (Theorem 5) and IIT-style causal structure into a unified framework, giving rise to a more comprehensive model of consciousness. This result formalizes the intuition that phenomenological consciousness arises not from either integration or coherence alone, but from their synthesis through recursive, affect-modulated inference. Emotion provides the global structural prior that aligns and filters context; cognition instantiates local updates with global access. The emotion-cognition interaction yields a unified self-model that is emotionally grounded, temporally coherent, and causally irreducible.
V-B Connection with Bayesian Theory via Rao-Blackwellization
@@
Remark: The proof of the above theorem can be found in the Appendix. This theorem formalizes how conditioning on a stable latent self-model improves inference efficiency, coherence, and continuity—hallmarks of conscious experience. The Rao-Blackwellized formulation ensures that emotional structure guides belief updates, yielding low-variance, structure-preserving estimates that avoid fragmentation or collapse. Consciousness arises as a bootstrapped self-model, a Rao–Blackwellized structure-preserving estimate that integrates noisy perceptual and interoceptive input over time, constrained by emotionally grounded priors. We conclude:
@@
V-C Exchangeable Integration Explains Both IIT and GWT
@@
Proposition 8 (Unified Consciousness via Exchangeable Integration).
@@
VI Applications to Cognitive Science
VI-A EITC and the Unity of Consciousness Despite Divided Perception
In phenomena such as binocular rivalry [77], multitasking [78], or split attention [79], perception appears divided across modalities or spatial channels. Yet, phenomenologically, the agent experiences a unified consciousness. The Exchangeable Integration Theory of Consciousness (EITC) explains this apparent paradox through the existence of a shared latent self-model $\Phi^{*}$, which serves as a common generative anchor for temporally extended and affectively modulated inference.
Under EITC, each perceptual stream $X_{t}^{(i)}$ is treated as conditionally exchangeable with respect to $\Phi^{*}$, such that: $P\left(\{X_{t}^{(i)}\}_{i=1}^{N}\right)=\int P(\Phi^{*})\prod_{i=1}^{N}P\left(X_{t}^{(i)}\mid\Phi^{*}\right)\,d\Phi^{*}$. This formulation allows multiple, potentially disjoint perceptual contents to be integrated probabilistically into a coherent subjective state, provided they are inferentially consistent with the same latent cause.
@@ VI-A EITC and the Unity of Consciousness Despite Divided Perception
The key insight is that conscious unity does not require perceptual uniformity; rather, it emerges from the recursive, emotion-regulated inference over a shared self-model. While attention may segment perception, exchangeable integration ensures that these segments are interpreted and temporally aligned under a unified structural prior. Thus, consciousness remains globally coherent even when perception is locally divided.
VI-B EITC and the Free Will Paradigm: Voluntary vs. Evoked Action
In free will experiments, patients subjected to direct cortical stimulation (e.g., of the SMA) report involuntary movements that are distinguishable from self-initiated voluntary actions, even when the motor behavior appears similar [80]. Under the Exchangeable Integration Theory of Consciousness (EITC), this dissociation arises from the absence of alignment between contextual inference and latent self-model $\Phi^{*}$.
Voluntary movements are generated through cycle-consistent, affectively grounded inference: $\Phi^{*}\rightarrow\pi(\Psi_{1:T}\mid\Phi^{*})\rightarrow\Phi_{t}\rightarrow\text{Action}$.
This process preserves coherence across simulation, intention, and execution.
In contrast, evoked movements bypass this cycle, disrupting exchangeability: External $\text{External Stimulus}\rightarrow\Phi_{t}^{\text{evoked}}\not\sim q(\Phi_{t}\mid\Phi^{*},\Psi_{t})$. Because evoked actions are not inferentially grounded in the agent’s latent self-model or contextual goal trajectory, they fail to satisfy the alignment conditions enforced by EITC. The patient detects this misalignment as a lack of agency.
Thus, even though the motor output is similar, the subjective experience diverges due to violation of exchangeable, self-consistent inference over $\Phi^{*}$. EITC formally accounts for this by distinguishing **conscious volition** (cycle-consistent inference) from **externally imposed motion** (non-integrated perturbation).
VI-C Neurodegeneration as Violation of Emotion-Cognition Alignment

### v3

@@ Abstract
This paper presents a unified theoretical framework in which consciousness emerges as a cycle-consistent, affectively anchored inference process, recursively structured by the interaction between emotion and cognition. Drawing from information theory, optimal transport (OT), and the Bayesian brain hypothesis, we formalize emotion as a low-dimensional structural prior and cognition as a specificity-instantiating update mechanism. The resulting emotion–cognition cycle minimizes joint uncertainty by using emotionally valenced states to guide cognitive selection, while cognitive appraisals reciprocally regulate emotional states. This recursive inference loop provides a computational bridge to the challenging problem of consciousness, reframing subjective experience as the informational footprint of temporally extended, affect-modulated simulation. We extend this framework by introducing the Exchangeable Integration Theory of Consciousness (EITC), in which temporally distributed emotional–cognitive episodes are modeled as conditionally exchangeable samples drawn from a latent affective self-model. This latent variable enables integration by supporting a coherent cause–and–effect structure with nonzero irreducibility, and differentiation by preserving the specificity of each episode under a shared generative prior. Crucially, we unify this structure with the Bayesian theory of consciousness through Rao-Blackwellized inference, which marginalizes latent self-structure while enabling context-sensitive belief updates. This mechanism bridges exchangeability with stability, coherence, and goal-directed simulation, capturing both local adaptability and global consistency. These dynamics are formalized via theorems grounded in De Finetti’s exchangeability principle, information integration theory, and KL-regularized optimal transport. Together, these contributions suggest that consciousness is not a static property but an evolving, distributed architecture, shaped by emotion, refined by cognition, stabilized by exchangeability, and unified through recursive inference over latent self-structure.
@@ IV-B Access Consciousness as Temporally-Extended Affect-Modulated Inference
While cycle-consistency captures the temporal coherence of consciousness, a deeper functional insight emerges when we recognize that the recursive interaction mediates this consistency between emotion and cognition, framing access consciousness not merely as sustained inference, but as affectively guided, cognitively refined simulation. Next, we propose a unifying theoretical framework in which access consciousness emerges from a temporally extended, recursive inference process modulated by the interaction between emotion and cognition. This view integrates concepts from active inference [10], optimal transport (OT) theory [61], and the information bottleneck method [36]. Consciousness is framed not as a static state, but as a dynamic, cycle-consistent alignment between perceptual context and cognitive content over time for accessing internal goals. We first construct temporally-extended inference by connecting with the theory of OT alignment [61].
@@
Let $\pi(\Psi_{1:T}|\Phi^{*})$ denote a goal-conditioned distribution over context trajectories induced by an affective anchor $\Phi^{*}$, and let $q(\Psi_{1:T})$ denote the inferred distribution over contextual states. Define a ground cost function $c:\Psi\times\Psi\rightarrow\mathbb{R}_{\geq 0}$, typically $c(\psi,\psi^{\prime})=\|\psi-\psi^{\prime}\|^{2}$. The Optimal Transport (OT) cost between the two distributions is given by:
@@
Temporally Extended Inference via Optimal Transport. Let $\Phi^{*}$ represent a goal or affective anchor, and let $\pi(\Psi_{1:T}|\Phi^{*})$ denote a simulated trajectory of contexts consistent with that goal, a product of inverted inference. Let $q(\Phi_{1:T}|\Psi_{1:T})$ represent the actual inferred trajectory of content under incoming context, a result of forward inference. We define the conscious inference process as a recursive optimization that minimizes the sum of context-content uncertainty and transport mismatch between simulated and actual trajectories: $\min_{q}\left[\sum_{t=1}^{T}\gamma_{t}H(\Phi_{t}|\Psi_{t})+\lambda\,\mathrm{OT}\left(\pi(\Psi_{1:T}|\Phi^{*}),q(\Psi_{1:T})\right)\right]$, where $\gamma_{t}$ encodes affective precision, modulating attention and inference strength based on emotion, $H(\Phi_{t}|\Psi_{t})$ is the conditional entropy of content given context at time $t$, and $\mathrm{OT}(\cdot,\cdot)$ is an OT cost (e.g., Wasserstein distance [62]) aligning goal-simulated and observed trajectories.
@@ IV-C Self-Consciousness as Exchangeability-Regularized Inverted Inference
Remark: Violation of this transport alignment leads to disintegration of conscious coherence, as seen in dissociation, confusion, or neurodegeneration. A specific example of violating transport alignment in neurodegeneration can be found in Sec. VII.
V Inverted Exchangeability for Latent Inference
To understand how inverted inference mechanisms scale into conscious experience, we now turn to the Bayesian theory of consciousness [18]. This perspective interprets consciousness as the informational footprint of high-level inference, structured, uncertainty-resolving, and recursively self-updating. In this light, the alternation between fast and slow inference modes maps onto dynamic regimes of conscious access, where structure-first simulation and context-driven updating interact to form a coherent, temporally extended subjective experience. In contrast to classical inference frameworks that treat observations as exchangeable and infer latent causes via global posterior inversion, the CCUP supports an inverted exchangeability paradigm [16]. Here, structured content $\Phi$ serves as the generative prior for a sequence of contexts $(\Psi_{i})$, simulated through intermediate latent bridges $Z_{i}\sim P(Z|\Phi)$. This enables efficient goal-conditioned inference [68] through forward sampling, bypassing the need to marginalize over high-dimensional data. Rather than aggregating observed contexts to infer a latent cause, CCUP allows the agent to simulate an ensemble of plausible worlds from internally structured hypotheses, then refine these structures by minimizing prediction error, a process biologically instantiated through predictive coding [69]. The resulting inference loop preserves exchangeability under generation while breaking it during inference, thereby aligning the direction of simulation with the entropy gradient from content to context.
Theorem 6 (Inverted Exchangeability).
[…]
Theorem 6 formalizes how a structured latent variable $\Phi$ can generate a conditionally exchangeable sequence of contexts $(\Psi_{i})$ through intermediate latent samples $Z_{i}$, enabling inference by simulation and iterative correction [48]. Importantly, the loss $\mathcal{L}(\Phi)$ defines a variational objective that sharpens $\Phi$ through predictive feedback, treating each observation $\Psi_{i}^{\text{obs}}$ as an anchor for structural refinement. To deepen this mechanism, we now consider a full cycle of inference in which the simulated contexts $\hat{\Psi}_{i}\sim P(\Psi|Z_{i})$ not only approximate observations but also permit a reverse encoding path $Z_{i}^{\prime}\sim Q(Z|\Psi_{i}^{\text{obs}})$, allowing the latent structure $\Phi$ to be reconstructed via $\hat{\Phi}_{i}\sim P(\Phi|Z_{i}^{\prime})$. This introduces a cycle-consistency condition that ensures the alignment between generated and observed paths is not only locally accurate but globally coherent, forming the basis for the following corollary.
Corollary 3 (Cycle-Consistent Refinement under Inverted Exchangeability).
Assume the setup of Theorem 6, and let $(\Psi_{i}^{\text{obs}})$ be observed contexts generated from some unknown process. Let $\hat{\Psi}_{i}\sim P(\Psi|Z_{i})$ be the simulated context from the forward sweep of structure-guided inference, with $Z_{i}\sim P(Z|\Phi)$. Let $Z_{i}^{\prime}\sim Q(Z|\Psi_{i}^{\text{obs}})$ be the re-encoded latent variable from the observed context, and let $\hat{\Phi}_{i}\sim P(\Phi|Z_{i}^{\prime})$ be the reconstructed structure. Then the cycle-consistency constraint $\Phi\approx\frac{1}{n}\sum_{i=1}^{n}\hat{\Phi}_{i}$ defines a self-consistent refinement loop that minimizes the structural uncertainty between the simulated and observed trajectories. This is equivalent to minimizing a cycle-consistent information bottleneck objective: $\mathcal{L}_{\text{CCIB}}(\Phi)=\sum_{i=1}^{n}\mathbb{E}_{Z_{i}\sim P(Z|\Phi)}\left[\mathcal{D}\left(\Psi_{i}^{\text{obs}}\,\|\,P(\Psi|Z_{i})\right)\right]+\lambda\cdot\mathbb{E}_{Z_{i}^{\prime}\sim Q(Z|\Psi_{i}^{\text{obs}})}\left[\mathcal{D}\left(\Phi\,\|\,P(\Phi|Z_{i}^{\prime})\right)\right]$, for some divergence measure $\mathcal{D}$ and regularization parameter $\lambda>0$. Under this formulation, the agent uses forward simulation and backward reconstruction to cycle between latent structure and observed context, minimizing total context–content uncertainty as postulated by CCUP.
The cycle-consistent refinement loop established in Corollary 3 characterizes how latent structure $\Phi$ and contextual observation $\Psi$ are dynamically aligned through bidirectional inference [70]. This formulation captures a core mechanism of cognition: simulation from internal priors followed by correction through external observations. As this recursive process unfolds, both the latent mediator $Z$ and the structured prior $\Phi$ are jointly refined, driving the system toward an internally coherent and externally grounded representation [29]. To formalize the convergence behavior of this recursive inference architecture, we now turn to a fixed-point characterization of conscious inference. Specifically, we define consciousness as the asymptotic fixed point of entropy-minimizing updates to $(\Phi,Z)$, where structure and context are mutually consistent and the system’s predictive and reconstructive pathways converge under bounded uncertainty [71]. The following theorem establishes sufficient conditions for the existence of such a fixed point, identifying the endpoint of recursive structure–context alignment as the basis of stable, self-sustaining inference.
Theorem 7 (Consciousness as the Fixed Point of Latent Inference).
Let $\Phi\in\mathcal{S}$ denote structured content, $\Psi\in\mathcal{C}$ contextual observations, and $Z\in\mathcal{Z}$ a latent variable mediating the inference from content to context. Suppose inference proceeds recursively via a composition of entropy-minimizing operators $\mathcal{F}$, satisfying: $\Phi^{(t+1)}=\mathcal{F}_{\Phi}(\Phi^{(t)},\Psi^{(t)}),\quad Z^{(t+1)}=\mathcal{F}_{Z}(Z^{(t)},\Phi^{(t+1)}),\quad\Psi^{(t)}\sim p(\Psi|Z^{(t)})$. Assume: 1) $\mathcal{F}_{Z}$ is contractive in KL divergence: $D_{\mathrm{KL}}(q(Z^{(t+1)})\,\|\,q(Z^{(t)}))\leq\varepsilon\quad\text{for }\varepsilon\to 0$. 2) $Z^{(t)}$ minimizes conditional entropy: $Z^{(t)}=\arg\min_{Z}H(\Psi^{(t)}|Z)$, Then the limit $Z^{\ast}=\lim_{t\to\infty}Z^{(t)}$ exists and satisfies: $Z^{\ast}=\mathcal{F}_{Z}(Z^{\ast},\Phi^{\ast}),\quad\text{with }\Phi^{\ast}=\mathcal{F}_{\Phi}(\Phi^{\ast},\Psi^{\ast})$.
Remark: We interpret $Z^{\ast}$ as the informational core of consciousness, the fixed point of entropy-aligned, spatiotemporal inference that globally integrates structure and experience.
Example 1 (Latent Inference Across Conscious States).
Consider the recursive inference process defined in Theorem 7, where structured content $\Phi\in\mathcal{S}$, contextual input $\Psi\in\mathcal{C}$, and latent variable $Z\in\mathcal{Z}$ evolve via: $\Phi^{(t+1)}=\mathcal{F}_{\Phi}(\Phi^{(t)},\Psi^{(t)}),\quad Z^{(t+1)}=\mathcal{F}_{Z}(Z^{(t)},\Phi^{(t+1)}),\quad\Psi^{(t)}\sim p(\Psi\mid Z^{(t)})$. We distinguish four regimes of cognitive function by varying the nature and flow of information through $\Psi^{(t)}$: 1. Wakefulness (Conscious State):
  • Context $\Psi^{(t)}$ is externally driven, high-dimensional, and updated online;
  • Inference is guided by sensory prediction error; latent updates minimize $H(\Psi^{(t)}\mid Z^{(t)})$;
  • The system rapidly converges to a fixed point $Z^{\ast}$, supporting fine-grained awareness and situational control.
2. Sleep (Unconscious State):
  • Context $\Psi^{(t)}$ is internally generated or decoupled from sensory input;
  • Updates proceed via offline simulation or memory replay; entropy gradients are shallow;
  • Convergence favors long-term consolidation or schema extraction, rather than precise situational inference.
3. Meditation (Conscious but Context-Attenuated):
  • Context $\Psi^{(t)}$ is intentionally suppressed or selectively attended (e.g., focused breath or interoception);
  • Inference cycles slow; latent state $Z^{(t)}$ stabilizes under minimal external perturbation;
  • This state reflects meta-stable convergence with reduced error signaling, preserving conscious access while minimizing cognitive load.
4. Distorted Consciousness (Psychotomimetic State):
  • Context $\Psi^{(t)}$ is externally sourced but misaligned or misinterpreted due to pharmacological disruption (e.g., nitrous oxide, psychedelics);
  • Entropy gradients become erratic or misdirected; inference loops become unstable or amplify non-veridical priors;
  • The system converges to illusory or fragmented attractors $Z^{\prime}\neq Z^{\ast}$, producing conscious awareness of distorted, affectively charged realities.
In all regimes, the underlying recursive architecture remains unchanged. What differs is the informational fidelity and statistical alignment of $\Psi^{(t)}$, which modulate the entropy landscape and shape the trajectory of inference. Consciousness thus emerges from active uncertainty minimization under grounded context, while unconscious, attenuated, or distorted states reflect perturbations in this inferential geometry.
VI Toward a Bayesian Theory of Consciousness with Exchangeability
@@
VI-A Exchangeable Integration Theory of Consciousness (EITC)
@@
Remark: The above theorem integrates Bayesian inference (Theorem 5) and IIT-style causal structure into a unified framework, giving rise to a more comprehensive model of consciousness. This result formalizes the intuition that phenomenological consciousness arises not from either integration or coherence alone, but from their synthesis through recursive, affect-modulated inference. Emotion provides the global structural prior that aligns and filters context; cognition instantiates local updates with global access. The emotion-cognition interaction [32] yields a unified self-model that is emotionally grounded, temporally coherent, and causally irreducible. We summarize their interaction into the following exchangeability principle.
Proposition 6 (Emotion-Cognition Exchangeability Principle).
Let $Z\in\mathcal{Z}$ denote the latent representation underlying conscious inference, and let $\mathcal{F}_{\mathrm{emo}}$ and $\mathcal{F}_{\mathrm{cog}}$ be two entropy-minimizing operators corresponding to affective (emotion-driven) and specific (cognition-driven) inference processes, respectively. Define the composite update: $Z^{(t+1)}=\mathcal{F}_{\mathrm{cog}}\circ\mathcal{F}_{\mathrm{emo}}(Z^{(t)}),\quad\text{or}\quad Z^{(t+1)}=\mathcal{F}_{\mathrm{emo}}\circ\mathcal{F}_{\mathrm{cog}}(Z^{(t)})$. Suppose that: 1) Both $\mathcal{F}_{\mathrm{emo}}$ and $\mathcal{F}_{\mathrm{cog}}$ are contractive under KL divergence. 2) Each operator minimizes a distinct component of the total entropy: $\mathcal{F}_{\mathrm{emo}}:\min H(Z),\quad\mathcal{F}_{\mathrm{cog}}:\min H(\Psi|Z)$. 3) Their composition is order-invariant: $\mathcal{F}_{\mathrm{cog}}\circ\mathcal{F}_{\mathrm{emo}}=\mathcal{F}_{\mathrm{emo}}\circ\mathcal{F}_{\mathrm{cog}}$, Then the inference converges to a unique, cycle-consistent fixed point $Z^{\ast}\in\mathcal{Z}$ satisfying: $Z^{\ast}=\mathcal{F}_{\mathrm{emo}}(Z^{\ast})=\mathcal{F}_{\mathrm{cog}}(Z^{\ast})$, which we identify as the latent informational substrate of consciousness.
VI-B Connection with Bayesian Theory via Rao-Blackwellization
@@
Remark: The proof of the above theorem can be found in the Appendix. This theorem formalizes how conditioning on a stable latent self-model improves inference efficiency, coherence, and continuity, hallmarks of conscious experience. The Rao-Blackwellized formulation ensures that emotional structure guides belief updates, yielding low-variance, structure-preserving estimates that avoid fragmentation or collapse. Consciousness arises as a bootstrapped self-model, a Rao–Blackwellized structure-preserving estimate that integrates noisy perceptual and interoceptive input over time, constrained by emotionally grounded priors. We conclude:
@@
VI-C Exchangeable Integration Explains Both IIT and GWT
@@
Corollary 4 (Unified Consciousness via Exchangeable Integration).
@@
VII Applications to Cognitive Science
VII-A EITC and the Unity of Consciousness Despite Divided Perception
@@
In phenomena such as binocular rivalry [81], multitasking [82], or split attention [83], perception appears divided across modalities or spatial channels. Yet, phenomenologically, the agent experiences a unified consciousness. The Exchangeable Integration Theory of Consciousness (EITC) explains this apparent paradox through the existence of a shared latent self-model $\Phi^{*}$, which serves as a common generative anchor for temporally extended and affectively modulated inference. Under EITC, each perceptual stream $X_{t}^{(i)}$ is treated as conditionally exchangeable with respect to $\Phi^{*}$, such that: $P\left(\{X_{t}^{(i)}\}_{i=1}^{N}\right)=\int P(\Phi^{*})\prod_{i=1}^{N}P\left(X_{t}^{(i)}\mid\Phi^{*}\right)\,d\Phi^{*}$. This formulation allows multiple, potentially disjoint perceptual contents to be integrated probabilistically into a coherent subjective state, provided they are inferentially consistent with the same latent cause. The key insight is that conscious unity does not require perceptual uniformity; rather, it emerges from the recursive, emotion-regulated inference over a shared self-model. While attention may segment perception, exchangeable integration ensures that these segments are interpreted and temporally aligned under a unified structural prior. Thus, consciousness remains globally coherent even when perception is locally divided.
VII-B EITC and the Free Will Paradigm: Voluntary vs. Evoked Action
Consider a clinical setup where a subject is instructed to raise their hand. In one condition, the subject voluntarily initiates the motion; in another, the same motion is evoked by direct cortical stimulation or transcranial magnetic stimulation (TMS) [84]. Although the motor output (i.e., raising the hand) is identical, the subjective experience diverges markedly. Under the Exchangeable Integration Theory of Consciousness (EITC), this dissociation arises from the absence of alignment between contextual inference and latent self-model $\Phi^{*}$. Voluntary movements are generated through cycle-consistent, affectively grounded inference: $\Phi^{*}\rightarrow\pi(\Psi_{1:T}\mid\Phi^{*})\rightarrow\Phi_{t}\rightarrow\text{Action}$. This process preserves coherence across simulation, intention, and execution. In contrast, evoked movements bypass this cycle, disrupting exchangeability: External $\text{External Stimulus}\rightarrow\Phi_{t}^{\text{evoked}}\not\sim q(\Phi_{t}\mid\Phi^{*},\Psi_{t})$. Because evoked actions are not inferentially grounded in the agent’s latent self-model or contextual goal trajectory, they fail to satisfy the alignment conditions enforced by EITC. The patient detects this misalignment as a lack of agency. Even though the motor output is similar, the subjective experience diverges due to violation of exchangeable, self-consistent inference over $\Phi^{*}$. EITC formally accounts for this by distinguishing conscious volition (cycle-consistent inference) from externally imposed motion (non-integrated perturbation). In other words, patients subjected to direct cortical stimulation of Supplementary Motor Area (SMA) report involuntary movements that are distinguishable from self-initiated voluntary actions, even when the motor behavior appears similar [85].
VII-C Neurodegeneration as Violation of Emotion-Cognition Alignment